the probability that a student takes calculus or is on scholarship is 0.85. The probability that a student is on scholarship is 0.61 and the probability that a student is taking calculus is 0.31. If a student is randomly chosen, find the probability that the student is taking calculus and is on scholarship.
Answer by ikleyn(52847)

.

Use the general formula of the probability theory


    P(A or B) = P(A) + P(B) - P(A and B).


Substitute the given values to get


    0.85 = 0.61 + 0.31 - P(calculus AND scholarship).


It implies


    P(calculus AND scholarship) = 0.61 + 0.31 - 0.85 = 0.07 = 7%.    ANSWER

Solved and carefully explained.